# FIMP Dark Matter in bulk viscous non-standard cosmologies

### v2

@@ II Non-standard cosmologies
As discussed in Refs. [107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120], a straightforward approach to generate NSC scenarios is adding in the early universe a field ($\phi$) which eventually decays into SM plasma. The evolution of the energy density, $\rho_{\phi}$, for this field is given by
@@
where “dot” denotes the derivative with respect to the cosmic time $t$, $\Gamma_{\phi}$ is the decay rate of $\phi$, $\omega\equiv p_{\phi}/\rho_{\phi}$ is the barotropic index with $p_{\phi}$ the pressure of the field, and $H=(\rho_{\phi}+\rho_{\gamma})/(3M_{p}^{2})$ is the Hubble parameter with $M_{p}=2.48\times 10^{18}$ GeV the reduced Planck mass. The SM bath evolution is given by
@@
The SM bath temperature is related with the radiation energy density by $\rho_{\gamma}=\pi^{2}g_{\star}(T)T^{4}/30$, with $g_{\star}$ the degrees of freedom that contribute to this energy density. It is important to mention that this NSC scenario must not alter the Big Bang Nucleosynthesis (BBN) epoch due to the highly accurate observations based on the $\Lambda$CDM model [121, 122, 123]. Therefore, the $\phi$ field must decay before the beginning of BBN, i.e, at $T_{\text{end}}\gtrsim T_{BBN}\sim 4$ MeV. Also, the temperature $T_{\text{end}}$ is defined as the moment when $H=\Gamma_{\phi}$ and can be related to the decay rate as
@@
It is important to mention that the addition of the field can influence the inflation and reheating epochs, as studied in various works [124, 125, 126, 127, 119, 128, 129]. While some propose that $\phi$ acts as the inflationary field within the NSC framework, others suggest that $\phi$ solely generates the reheating epoch [130, 131, 111, 115].
@@
II.1 FIMPs in non-standard cosmologies
FIMPs are generated by a non-thermal production mechanism called Freeze-In. This DM genesis is based on the assumption that if the coupling of DM particles to the SM bath is small, meaning that their interactions with these particles are very feebly, then the DM particle never achieves the thermal equilibrium in the early universe with the bath particles. With the ongoing expansion of the universe, the production of FIMPs come to an end as their interaction rates become increasingly inefficient relative to the expansion rate of the universe, leading to the observed DM abundance.
These candidates can establish their abundance through two distinct mechanisms. The first one is UltraViolet (UV) freeze-in, and the second one is InfraRed (IR) freeze-in. In the UV freeze-in scenario, the process occurs when the mass of the mediator particle of the interaction with FIMPs exceeds the temperature of the thermal bath. This interaction involves higher-order, non-renormalizable operators, making it highly sensitive to the reheating temperature in the early universe. On the other hand, the IR freeze-in mechanism takes place at temperatures comparable to the mass of DM, through renormalizable operators. For example, in Ref. [132] it is explore DM UV freeze-in from a thermal bath during inflation (a warm inflation phase), showing that the DM yield has a non-trivial dependence on the temperature and the Hubble rate during inflation.
@@ II.1 FIMPs in non-standard cosmologies
with $\Lambda$ an energy dimension quantity related to the mass scale of the DM and SM bath mediator. Following this line, operators of mass dimension 5 ($n=0$) give constant values of $\langle\sigma v\rangle$, with some examples presented in Refs. [133, 134, 22, 135, 136, 137]. Meanwhile, examples of operators with mass dimensions 6 ($n=2$) or 7 ($n=4$) are shown in Refs. [138, 139, 134, 140]. Along this paper, we mainly analyze the case with a constant $\langle\sigma v\rangle$, and briefly discuss the cases with $n=2$ and $n=4$ in Subsec. III.3.
In the case of the NSC, the decay of the new field, introduced at early times, generates an increment in the temperature of the universe and, consequently, in the entropy and radiation energy density. This change in the temperature can also be seen as a faster or slower expansion rate of the universe. The entropy injection can be parametrized by $D\equiv s(T_{\text{end}})/s(m_{\chi})=\left(T_{\text{end}}/m_{\chi}\right)^{3}$, i.e, the entropy density before and after the decays of the new field $\phi$. This left significant imprints on the DM production, considering that the Yield depends explicitly on the entropy density. Therefore, an increment in $s(T)$ will dilute the DM relic density. The latter means that DM parameters ($m_{\chi}$, $\langle\sigma v\rangle$) that overproduce the relic density in the $\Lambda$CDM model can, in the NSC scenario, reproduce the current DM relic density. The establishment of DM can occur in the four regions described in Section II:
@@
Figure 3: Comparison in the yield production between the NSC (solid blue line) and the $\Lambda$CDM (dashed-dotted red line) scenarios for $\kappa=10^{-2}$, $T_{\text{end}}=7\times 10^{-3}$ GeV, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=5\times 10^{-26}$ GeV-2, and $\omega=0$. The dashed lines correspond to $x_{\text{eq}}$ (cyan), $x_{\text{c}}$ (grey), and $x_{\text{end}}$ (green); while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. [9], observing that the parameter space considered in this case gives us the right amount of DM in the NSC scenario.
@@
Figure 4: Comparison in the yield production between the NSC (solid blue line) and the $\Lambda$CDM (dashed-dotted red line) scenarios for $\kappa=1$, $T_{\text{end}}=5\times 10^{-2}$ GeV, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=10^{-24}$ GeV-2, and $\omega=0$. The dashed lines correspond to $x_{\text{c}}$ (grey) and $x_{\text{end}}$ (green), while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. [9], observing that the parameter space considered in this case gives us the right amount of DM in the NSC scenario.
@@
Figure 5: Comparison in the yield production between the NSC (solid blue line) and the $\Lambda$CDM (dashed-dotted red line) scenarios for $\kappa=10^{3}$, $T_{\text{end}}=2$ GeV, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=5\times 10^{-22}$ GeV-2, and $\omega=0$. The dashed lines correspond to $x_{\text{c}}$ (grey) and $x_{\text{end}}$ (green), while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. [9], observing that the parameter space considered in this case gives us the right amount of DM in the NSC scenario.
@@
It is important to note that in the case of UV Freeze-in, the DM production can also be generated by other mechanisms, such as the decay of $\phi$ or gravitational production. According to Ref. [117], considering couplings of $\mathcal{O}(1)$, $\omega\sim 1$, and $m_{\phi}<M_{p}$, the contribution of the Branching ratio of $\phi$ decaying in DM is negligible compared to the UV Freeze-in. In the case of gravitational production, the interaction rate density is proportional to $T^{8}/M_{p}^{4}$, meaning that higher initial temperatures ($T_{\text{ini}}>10^{11}$ GeV) and higher reheating temperatures $(T_{\text{rh}}>10^{7}$ GeV) would provide a perfect mechanism for generating FIMP DM [141, 142]. In the case we are analyzing, the form of $\langle\sigma v\rangle$ is more efficient.
III Bulk viscous non-standard cosmologies
@@
In the bulk viscous extension of the classical NSC, it is assumed that the field $\phi$ experiences a dissipative process in the form of a bulk viscosity during its cosmic evolution. In this extension, and working in the relativistic thermodynamic framework out of equilibrium of Eckart’s theory [50], the equations for the Hubble parameter and the evolution of the SM bath (2) remain unchanged, while the acceleration equation and conservation equation for the field (1) become
@@
where $P_{\text{eff},\phi}=p_{\phi}+\Pi$ is the effective pressure of the field, with $p_{\phi}$ being the equilibrium pressure, $\Pi=-3H\xi$ the bulk viscous pressure, and $\xi$ is the bulk viscosity. In particular, we assume that the dissipative fluid does not experience heat flow or shear viscosity. For more details on the procedure used to obtain this novel NSC, we refer the reader to Ref. [106], where the Bulk viscous NSCs were studied for the first time. It is important to note that the bulk viscosity influences the evolution of the universe through the bulk viscous pressure. In particular, in an expanding universe, the expression $\Pi=-3H\xi$ is always negative ($\xi>0$ to maintain consistency with the second law of thermodynamics [143]). Consequently, this viscosity leads to an acceleration in the expansion of the universe, as seen from Eq. (7).
In general, bulk viscosity can depend on the temperature and pressure of the dissipative fluid [143]. Hence, a natural and most extensively studied expression for the bulk viscosity is to assume a proportional dependence to the power of its energy density, $\xi=\xi_{0}\rho_{\phi}^{1/2}$, where $\xi_{0}=\hat{\xi_{0}}M_{p}$, allowing $\hat{\xi_{0}}$ to be a dimensionless parameter. With this consideration, Eq. (8) takes the form
@@
III.1 Comparison between scenarios
In this section, we will compare all the features discussed in Sections II and III, between the NSC and the bulk viscous NSC scenarios.
In Fig. 6, we depict the evolution of $\rho\times(a/a_{0})^{4}$ as a function of the temperature $T$ for $\kappa=10^{-3}$, $T_{\text{end}}=7\times 10^{-3}$ GeV, $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, and $\omega=0$. The solid and dashed-dotted lines correspond to the NSC with and without bulk viscosity, respectively; while the red lines represent the new field $\phi$ and the blue lines the radiation component. We also depict the values of $T_{\text{eq}}$ (cyan), $T_{\text{c}}$ (grey), and $T_{\text{end}}$ (green) for the NSC with bulk viscosity (dashed lines) and the classical NSC (dotted lines). From the figure, it can be observed that the bulk viscous NSC leads to an enhanced production of the field $\phi$ compared to the classical NSC scenario, resulting in a greater increase in the radiation energy density due to the decay of the field, and consequently, a higher entropy injection into the SM bath. An interesting feature is that the overall temperature-dependent behavior of the fluids remains largely unchanged between the two scenarios, allowing us to conclude that the field $\phi$ becomes negligible in the bulk viscous NSC scenario as in the classical NSC. Nevertheless, this conclusion is only valid for bulk viscous NSC models with $\omega>-1$, since viscosity can lead to obtaining an effective barotropic index where $\omega_{\text{eff}}\leq-1$ for an $\omega>-1$ but close to $-1$, i.e., we can obtain a behavior where the field $\phi$ never decays. Meanwhile, Figure 7 shows the comparison in the yield production between the classical NSC (solid red line) and the NSC with bulk viscosity (solid blue line) for the same values of $\kappa$, $T_{\text{end}}$, $\hat{\xi_{0}}$, $m_{\chi}$, and $\omega$ as in Figure 6, considering $\langle\sigma v\rangle=5\times 10^{-26}$ GeV-2. The dashed and dotted lines correspond to $x_{\text{eq}}$ (cyan), $x_{\text{c}}$ (grey), and $x_{\text{end}}$ (green) for the NSC with and without bulk viscosity, respectively; while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. [9]. Firstly, note that the lines for $x_{\text{end}}$ and $x_{\text{fi}}$ are the same for both models. Secondly, note that the behavior of the Yield production is the same in both models until the decay of $\phi$. After that, it is observed that the higher entropy injection generated by the viscous term leads to lower values of the DM Yield. This means that the DM relic density agrees with the current observation for a NSC with bulk viscosity in contrast with the standard NSC case which overproduces the DM relic density.

### v3

@@ 2 Non-standard cosmologies
As discussed in Refs. Giudice et al. (2001); Hamdan and Unwin (2018); D’Eramo et al. (2018, 2017); Visinelli (2018); Drees and Hajkarim (2018); Bernal et al. (2019a, b); Maldonado and Unwin (2019); Arias et al. (2019); Bernal et al. (2019); Arias et al. (2021); Bernal and Xu (2022); Silva-Malpartida et al. (2024), a straightforward approach to generate NSC scenarios is adding in the early universe a field ($\phi$) which eventually decays into SM plasma. The evolution of the energy density, $\rho_{\phi}$, for this field is given by
@@
where “dot” denotes the derivative with respect to the cosmic time $t$, $\Gamma_{\phi}$ is the decay rate of $\phi$, $\omega\equiv p_{\phi}/\rho_{\phi}$ is the barotropic index with $p_{\phi}$ the pressure of the field, and $H=(\rho_{\phi}+\rho_{\gamma})/(3M_{p}^{2})$ is the Hubble parameter with $M_{p}=2.48\times 10^{18}$ GeV the reduced Planck mass. In the same way as for $\phi$, the equation for the radiation energy density, $\rho_{\gamma}$, can also be written. However, to be more precise, it is necessary to incorporate the relativistic degrees of freedom for radiation. Therefore, the evolution of the Standard Model (SM) bath is more accurately described by the entropy density, due to the entropy injection resulting from the decay of $\phi$ Drees and Hajkarim (2018). Consequently, the evolution of the SM bath is given by
@@
The SM bath temperature is related with the radiation energy density by $\rho_{\gamma}=\pi^{2}g_{\star}(T)T^{4}/30$, with $g_{\star}$ the degrees of freedom that contribute to this energy density. It is important to mention that this NSC scenario must not alter the Big Bang Nucleosynthesis (BBN) epoch due to the highly accurate observations based on the $\Lambda$CDM model Sarkar (1996); Hannestad (2004); De Bernardis et al. (2008). Therefore, the $\phi$ field must decay before the beginning of BBN, i.e, at $T_{\text{end}}\gtrsim T_{BBN}\sim 4$ MeV. Also, the temperature $T_{\text{end}}$ is defined as the moment when $H=\Gamma_{\phi}$ and can be related to the decay rate as
@@
It is important to mention that the addition of the field can influence the inflation and reheating epochs, as studied in various works Kofman et al. (1997); Ackerman et al. (2011); Garcia et al. (2020, 2021); Bernal and Xu (2022); Silva-Malpartida et al. (2023); Haque et al. (2024). While some propose that $\phi$ acts as the inflationary field within the NSC framework, others suggest that $\phi$ solely generates the reheating epoch Gelmini and Gondolo (2006); Gelmini et al. (2006); Visinelli (2018); Maldonado and Unwin (2019).
@@
2.1 FIMPs in non-standard cosmologies
The Boltzmann equation described in eq. (5) is valid for both WIMP and FIMP candidates. The main differences between them are the magnitude of the thermally averaged cross-section and the initial conditions: WIMPs have an initial number density, as they were in equilibrium with the Standard Model (SM) bath in the early universe, whereas FIMPs are produced via decays of other particles in the early universe and therefore have an initial number density of zero. This production of FIMPs particles is generated by a non-thermal production mechanism called Freeze-In. This DM genesis is based on the assumption that if the coupling of DM particles to the SM bath is small, meaning that their interactions with these particles are very feebly, then the DM particle never achieves the thermal equilibrium in the early universe with the bath particles. With the ongoing expansion of the universe, the production of FIMPs come to an end as their interaction rates become increasingly inefficient relative to the expansion rate of the universe, leading to the observed DM abundance.
These candidates can establish their abundance through two distinct mechanisms. The first one is UltraViolet (UV) freeze-in, and the second one is InfraRed (IR) freeze-in. In the UV freeze-in scenario, the process occurs when the mass of the mediator particle of the interaction with FIMPs exceeds the temperature of the thermal bath. This interaction involves higher-order, non-renormalizable operators, making it highly sensitive to the reheating temperature in the early universe. On the other hand, the IR freeze-in mechanism takes place at temperatures comparable to the mass of DM, through renormalizable operators. For example, in Ref. Freese et al. (2024) it is explore DM UV freeze-in from a thermal bath during inflation (a warm inflation phase), showing that the DM yield has a non-trivial dependence on the temperature and the Hubble rate during inflation.
@@ 2.1 FIMPs in non-standard cosmologies
with $\Lambda$ an energy dimension quantity related to the mass scale of the DM and SM bath mediator. Following this line, operators of mass dimension 5 ($n=0$) give constant values of $\langle\sigma v\rangle$, with some examples presented in Refs. Chowdhury et al. (2019); Elahi et al. (2015); Hall et al. (2010); Biswas et al. (2020); Shakya (2016); McDonald and Sahu (2009). Meanwhile, examples of operators with mass dimensions 6 ($n=2$) or 7 ($n=4$) are shown in Refs. Lebedev et al. (2012); Bernal et al. (2018); Elahi et al. (2015); Krauss et al. (2014). Along this paper, we mainly analyze the case with a constant $\langle\sigma v\rangle$, and briefly discuss the cases with $n=2$ and $n=4$ in Subsec. 3.3.
In the case of the NSC, the decay of the new field, introduced at early times, generates an increment in the temperature of the universe and, consequently, in the entropy and radiation energy density. This change in the temperature can also be seen as a faster or slower expansion rate of the universe. The entropy injection can be parametrized by $D\equiv s(T_{\text{end}})/s(m_{\chi})=\left(T_{\text{end}}/m_{\chi}\right)^{3}$, i.e, the entropy density before and after the decays of the new field $\phi$. This left significant imprints on the DM production, considering that the Yield depends explicitly on the entropy density. Therefore, an increment in $s(T)$ will dilute the DM relic density. The latter means that DM parameters ($m_{\chi}$, $\langle\sigma v\rangle$) that overproduce the relic density in the $\Lambda$CDM model can, in the NSC scenario, reproduce the current DM relic density. The establishment of DM can occur in the four regions described in Section 2:
@@
Figure 3: Comparison in the yield production between the NSC (solid blue line) and the $\Lambda$CDM (dashed-dotted red line) scenarios for $\kappa=10^{-2}$, $T_{\text{end}}=7\times 10^{-3}$ GeV, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=5\times 10^{-26}$ GeV-2, and $\omega=0$. The dashed lines correspond to $x_{\text{eq}}$ (cyan), $x_{\text{c}}$ (grey), and $x_{\text{end}}$ (green); while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. Aghanim et al. (2020), observing that the parameter space considered in this case gives us the right amount of DM in the NSC scenario.
@@
Figure 4: Comparison in the yield production between the NSC (solid blue line) and the $\Lambda$CDM (dashed-dotted red line) scenarios for $\kappa=1$, $T_{\text{end}}=5\times 10^{-2}$ GeV, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=10^{-24}$ GeV-2, and $\omega=0$. The dashed lines correspond to $x_{\text{c}}$ (grey) and $x_{\text{end}}$ (green), while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. Aghanim et al. (2020), observing that the parameter space considered in this case gives us the right amount of DM in the NSC scenario.
@@
Figure 5: Comparison in the yield production between the NSC (solid blue line) and the $\Lambda$CDM (dashed-dotted red line) scenarios for $\kappa=10^{3}$, $T_{\text{end}}=2$ GeV, $m_{\chi}=100$ GeV, $\langle\sigma v\rangle=5\times 10^{-22}$ GeV-2, and $\omega=0$. The dashed lines correspond to $x_{\text{c}}$ (grey) and $x_{\text{end}}$ (green), while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. Aghanim et al. (2020), observing that the parameter space considered in this case gives us the right amount of DM in the NSC scenario.
@@
It is important to note that in the case of UV Freeze-in, the DM production can also be generated by other mechanisms, such as the decay of $\phi$ or gravitational production. According to Ref. Bernal et al. (2019), considering couplings of $\mathcal{O}(1)$, $\omega\sim 1$, and $m_{\phi}<M_{p}$, the contribution of the Branching ratio of $\phi$ decaying in DM is negligible compared to the UV Freeze-in. In the case of gravitational production, the interaction rate density is proportional to $T^{8}/M_{p}^{4}$, meaning that higher initial temperatures ($T_{\text{ini}}>10^{11}$ GeV) and higher reheating temperatures $(T_{\text{rh}}>10^{7}$ GeV) would provide a perfect mechanism for generating FIMP DM Barman et al. (2022); Bernal et al. (2025). In the case we are analyzing, the form of $\langle\sigma v\rangle$ is more efficient.
3 Bulk viscous non-standard cosmologies
@@
In the bulk viscous extension of the classical NSC, it is assumed that the field $\phi$ experiences a dissipative process in the form of a bulk viscosity during its cosmic evolution. In this extension, and working in the relativistic thermodynamic framework out of equilibrium of Eckart’s theory Eckart (1940b), the equations for the Hubble parameter and the evolution of the SM bath (2) remain unchanged, while the acceleration equation and conservation equation for the field (1) become
@@
where $P_{\text{eff},\phi}=p_{\phi}+\Pi$ is the effective pressure of the field, with $p_{\phi}$ being the equilibrium pressure, $\Pi=-3H\xi$ the bulk viscous pressure, and $\xi$ is the bulk viscosity. In particular, we assume that the dissipative fluid does not experience heat flow or shear viscosity. For more details on the procedure used to obtain this novel NSC, we refer the reader to Ref. González et al. (2024), where the Bulk viscous NSCs were studied for the first time. It is important to note that the bulk viscosity influences the evolution of the universe through the bulk viscous pressure. In particular, in an expanding universe, the expression $\Pi=-3H\xi$ is always negative ($\xi>0$ to maintain consistency with the second law of thermodynamics Weinberg and Dicke ). Consequently, this viscosity leads to an acceleration in the expansion of the universe, as seen from Eq. (7).
In general, bulk viscosity can depend on the temperature and pressure of the dissipative fluid Weinberg and Dicke . Hence, a natural and most extensively studied expression for the bulk viscosity is to assume a proportional dependence to the power of its energy density, $\xi=\xi_{0}\rho_{\phi}^{1/2}$, where $\xi_{0}=\hat{\xi_{0}}M_{p}$, allowing $\hat{\xi_{0}}$ to be a dimensionless parameter. With this consideration, Eq. (8) takes the form
@@
3.1 Comparison between scenarios
In this section, we will compare all the features discussed in Sections 2 and 3, between the NSC and the bulk viscous NSC scenarios.
In Fig. 6, we depict the evolution of $\rho\times(a/a_{0})^{4}$ as a function of the temperature $T$ for $\kappa=10^{-3}$, $T_{\text{end}}=7\times 10^{-3}$ GeV, $\hat{\xi_{0}}=10^{-2}$, $m_{\chi}=100$ GeV, and $\omega=0$. The solid and dashed-dotted lines correspond to the NSC with and without bulk viscosity, respectively; while the red lines represent the new field $\phi$ and the blue lines the radiation component. We also depict the values of $T_{\text{eq}}$ (cyan), $T_{\text{c}}$ (grey), and $T_{\text{end}}$ (green) for the NSC with bulk viscosity (dashed lines) and the classical NSC (dotted lines). From the figure, it can be observed that the bulk viscous NSC leads to an enhanced production of the field $\phi$ compared to the classical NSC scenario, resulting in a greater increase in the radiation energy density due to the decay of the field, and consequently, a higher entropy injection into the SM bath. An interesting feature is that the overall temperature-dependent behavior of the fluids remains largely unchanged between the two scenarios, allowing us to conclude that the field $\phi$ becomes negligible in the bulk viscous NSC scenario as in the classical NSC. Nevertheless, this conclusion is only valid for bulk viscous NSC models with $\omega>-1$, since viscosity can lead to obtaining an effective barotropic index where $\omega_{\text{eff}}\leq-1$ for an $\omega>-1$ but close to $-1$, i.e., we can obtain a behavior where the field $\phi$ never decays. Meanwhile, Figure 7 shows the comparison in the yield production between the classical NSC (solid red line) and the NSC with bulk viscosity (solid blue line) for the same values of $\kappa$, $T_{\text{end}}$, $\hat{\xi_{0}}$, $m_{\chi}$, and $\omega$ as in Figure 6, considering $\langle\sigma v\rangle=5\times 10^{-26}$ GeV-2. The dashed and dotted lines correspond to $x_{\text{eq}}$ (cyan), $x_{\text{c}}$ (grey), and $x_{\text{end}}$ (green) for the NSC with and without bulk viscosity, respectively; while the dashed-dotted magenta line is the time when the DM candidate Freeze-In their number at $x_{\text{fi}}$. The green zone represents the current DM relic density according to Ref. Aghanim et al. (2020). Firstly, note that the lines for $x_{\text{end}}$ and $x_{\text{fi}}$ are the same for both models. Secondly, note that the behavior of the Yield production is the same in both models until the decay of $\phi$. After that, it is observed that the higher entropy injection generated by the viscous term leads to lower values of the DM Yield. This means that the DM relic density agrees with the current observation for a NSC with bulk viscosity in contrast with the standard NSC case which overproduces the DM relic density.